# Stratification governs the Existence of Surface-Intensified Eastward Jets in Turbulent Gyres without Bottom Friction

### v1

@@ Abstract
Eastward oceanic jets at mid-latitudes, such as the Gulf Stream or the Kuroshio, are important in oceanic circulation as they transport water masses from western boundary currents far into the open ocean. This study uses an idealized model to isolate the role of density stratification on the formation of such jets in turbulent gyres. Our findings describe density profiles that are favourable for jet formation and highlight the importance of strong stratification close to the surface.
@@ 1 Introduction
Surface-intensified eastward jets detaching from western boundaries are a prominent feature of mid-latitude oceans. Two iconic examples are the Gulf Stream and the Kuroshio, which appear as coherent, narrow, meandering ribbons surrounded by mesoscale oceanic rings in snapshots of surface kinetic energy (Chassignet and Xu, 2021). Satellite altimetry and state-estimate reconstructions show that these jets often maintain a coherent structure over long distances (Forget et al., 2015), despite the transition from coastal to open-ocean environments (Chassignet and Marshall, 2008). Coherent eastward jets are also found across a wide range of models, from high-resolution primitive equation models (Ajayi et al., 2020) to simple two-layer quasi-geostrophic models (Holland, 1978; Berloff and McWilliams, 1999a). In spite of its robustness, a simple theoretical description of the eastward jet in turbulent, western-intensified gyres is still lacking.
Eastward jets in quasi-geostrophic gyres have attracted significant attention over the last few decades. One approach, based on dynamical system theory, begins with viscous solutions (Stommel, 1948; Munk, 1950) and examines successive bifurcations as dissipation is reduced (Veronis, 1966; Cessi and Ierley, 1995; Dijkstra and Ghil, 2005). In this context, Simonnet (2005) described a weakly time-dependent eastward jet controlled by dissipative mechanisms in a one-layer quasi-geostrophic model. However, reaching sufficiently low dissipation regimes relevant to the real ocean remains a challenging task, as it leads to the excitation of an increasing number of degrees of freedom. A complementary theoretical approach has focused on constructing solutions for zones of inertial recirculation when the boundary layer approximation breaks down (Cessi et al., 1987; Ierley, 1990). Nakano et al. (2008) outlined a solution where an eastward jet flanked by inertial recirculation reconnects to a Sverdrup interior, but emphasize the difficulty of correctly describing the dynamical balance of the recirculation zones. Both their stability and driving mechanisms (Holland and Schmitz, 1985; Sun et al., 2013) pose a tenacious problem for the closure of mathematical models describing such inertial recirculation embedded in turbulent gyres (Kubokawa, 2023).
In the limit of vanishing viscosity, equilibrium statistical mechanics has been employed to predict the self-organization of ocean currents on basin scales. On the beta plane, at low energy, equilibrium statistical mechanics consistently yields Fofonoff flows (Salmon et al., 1976; Bretherton and Haidvogel, 1976). These solutions are characterized by strong jets, but these jets are westward and confined only to the boundaries. At sufficiently higher energy levels, 1.5-layer quasi-geostrophic models with a small Rossby radius of deformation predict the formation of intense zonal jets in the interior, separating regions of homogenized vorticity (Venaille and Bouchet, 2011). While these invicid approaches can lead to meaningful descriptions of inertial recirculation regions (Marshall and Nurser, 1986), they fail to account for western intensification when applied at the scale of an ocean basin. To wit, ocean gyres are strongly out of equilibrium turbulent systems.
A challenging aspect of mesoscale turbulence in ocean gyres is its propensity to rectify the large-scale flow that initiated small-scale motion through instabilities. In that context, baroclinic turbulence has often been adressed by looking at the equilibration of quasi-geostrophic fluctuations around a prescribed unstable mean flow in doubly periodic domains (Salmon, 1998; Arbic and Flierl, 2004b). This ”local” perspective has helped interpret observed flow patterns, such as vortices and meandering jets, based on the observed local mean flow in different regions of the ocean (Venaille et al., 2011), and it has also lead to the design of useful parameterizations of baroclinic turbulence in coarser models (Gallet and Ferrari, 2020; Hadjerci and Gallet, 2023, 2024). Among the variety of reported baroclinic turbulent regimes (Arbic and Flierl, 2004a), formation of coherent jets have been described (Arbic and Flierl, 2003; Venaille et al., 2014). While these patterns share similarities with Gulf Stream or Kuroshio jets they were only found in the limit of large bottom friction, which is not relevant for the ocean. Moreover, the locality assumption behind studies in doubly periodic geometries has been strongly questioned (Grooms et al., 2013). Thus, a minimal model for coherent eastward jets must account for non-local processes and move beyond the large friction assumption.
The main goal of the present work is to isolate the role of stratification in oceanic jet formation. Therefore we employ a quasi-geostrophic model of mid-latitude gyres and consider a scenario without bottom friction. Although bottom friction is needed to stabilize the energy cycle of unbounded geostrophic turbulence (Salmon, 1998), a significant amount of energy is supposed to be dissipated by flow-topography interactions close to the western boundary in more realistic settings (Dewar and Hogg, 2010; Molemaker et al., 2015; Deremble et al., 2017). We aim to reproduce this so-called ”western eddy-graveyard” (Zhai et al., 2010; Yang et al., 2021) by applying no-slip boundary conditions, which may act as an efficient energy sink while remaining in a quasi-geostrophic modelling framework (Miller et al., 2024).
@@
The structure of the article is as follows: In section 2, we describe the two-layer model and introduce the reduced parameter space which describes the vertical stratification properties. In section 3, we use numerical simulations to identify three distinct stratification regimes: strong, intermediate, and weak. We then shift our attention to the intermediate stratification regime. Section 4 presents a local linear stability analysis of surface-intensified gyres, paying particular attention to the distinct stability characteristics of the eastward and westward parts of the gyres, as well as the western boundary layer. In section 5, we explore the inviscid mechanisms behind the formation of zonally penetrating jets. We compare a two-layer and a one-and-a-half-layer quasi-geostrophic simulations to identify the dynamics specifically linked to baroclinic instability. Finally, in section 6, we discuss the transition to a zonostrophic regime, which precedes the weak stratification regime characterized by barotropic gyres. We conclude in section 7 by outlining the minimal conditions required for the emergence and persistence of a coherent jet between the gyres.
@@ 2 Flow Model
with $\psi_{i}$ the streamfunction and $\mathbf{u}_{i}=(u_{i},v_{i})=(-\partial_{y}\psi_{i},\partial_{x}\psi_{i})$ the velocity in each layer. We solve this system of equations on a square domain with $(x,y)\in[0,L]^{2}$, and apply no-slip boundary conditions.
with $g^{\prime}$ being the reduced gravity between both layers and $f_{0}$ the Coriolis parameter. In addition to $\delta$, the problem admits three non-dimensional parameters:
$\tilde{\nu}=\frac{\nu}{U_{Sv}L},\quad\tilde{\beta}=\frac{\beta L^{2}}{U_{Sv}},\quad\xi=\frac{U_{Sv}}{\beta L_{d}^{2}},\quad$ (6)
with
$U_{Sv}=\frac{8\pi\tau_{0}}{H_{1}L\beta}.$ (7)
The velocity $U_{Sv}$ gives the order of magnitude of the zonal flow in the gyres when circulation is confined in the upper layer and is set to $U_{Sv}=0.09$ ms-1 in all simulations. The parameters $\tilde{\nu}$ and $\tilde{\beta}$ are also kept constant throughout this study, such that the thickness of the inertial boundary layer is always $\delta_{I}=L/\sqrt{\tilde{\beta}}=72$ km (Charney (1955)) and the viscous sublayer is always $\delta_{P}=L\sqrt{\tilde{\nu}/\sqrt{\tilde{\beta}}}=3$ km (Pedlosky, 1987; Ierley and Ruehr, 1986) for a basin of size $L=4000$ km. We will focus on examining the dynamical regimes that appear upon varying the two parameters $\xi$ and $\delta$.
The parameter $\delta$ is a proxy for surface-intensified stratification, but small values of $\delta$ do not always imply surface-intensified currents as the system may adjust to settle into a barotropic state. The parameter $\xi$ is sometimes called the criticality parameter, related to baroclinic instability (Held and Larichev, 1996; Jansen and Ferrari, 2012). In order to reach asymptotic behaviour, we vary $\xi$ from $10^{-2}$ to $10^{2}$ for $\delta=0.5$, from $1$ to $5.6$ for $\delta=1/6$ and from $10^{-1}$ to $10^{3}$ for $\delta=0.01$. This corresponds to changing $L_{d}$ from $737$ km to $7$ km, $72$ km to $13$ km and $233$ km to $2$ km, respectively. In our numerical simulations, we change $L_{d}$ by setting $g^{\prime}$, which controls $\xi$ but leaves all other non-dimensional variables unchanged. A summary of the dimensional variables of the simulations is given in table 1.
To numerically solve equations (1) - (4) the code qgw was used. It uses classical finite differences discretization of the equations with an Arakawa formulation of the Jacobian and a direct spectral inversion (after projection on the vertical modes) to obtain the streamfunction $\psi$ from the potential vorticity $q$. Time stepping is done with an adaptive 2nd order Adams-Bashforth technique, where the time step is determined by a standard CFL-condition.
Unless specified, all runs are performed with 4096$\times$4096 grid points on the horizontal, corresponding to a resolution of $1$ km, ensuring all relevant flow scales of the problem are well resolved. We paid attention to run the simulations for at least two gyre turnover times $T_{Gyre}=L/U_{Sv}$ after the spinup was completed. Last, the no-slip boundary condition is implemented as in Miller et al. (2024).
@@
$50\times 10^{-5}$ m2 s-2,
Wind Stress $\tau_{0}$ $16\times 10^{-5}$ m2 s-2,
$1\times 10^{-5}$ m2 s-2
@@ 3.1 Reference Run
Figure 1: Upper layer (top row) and lower layer (bottom row) for the reference run at $\delta=0.01$, $\xi=10$. (A, D) Flow speed $|\mathbf{u}_{i}|$ normalized with inertial velocity $U_{I}=\psi_{Sv}/\delta_{I}$, (B, E) potential vorticity normalised by planetary vorticity and (C, F) time-mean streamfunction normalised by Sverdrup scaling.
@@ 3.2 Parameter Space
The focus of this manuscript lies on the intermediate stratification regime, due to the appearance of eastward jets in gyres such as presented in the reference run. This regime is characterised by $1<\xi<1/\delta$ and is hence accessible only when $\delta\ll 1$. We therefore focus on the runs at $\delta=0.01$, limiting our description to the upper layer as the dynamics remain mostly surface-intensified (figure 3). The surface flow pattern are displayed in figure 4. As $\xi$ is increased, a continuous transition occurs from western-intensified gyres to a state that restores east-west symmetry, with a decrease in total transport of the gyres. Coincidentally, the dominant turbulent flow features change. The strong coherent vortices which dominate the flow in the strong stratification regime disappear, and instead a sea of eddies of characteristic size $L_{d}$ develops in the regions of westward flow. In the regions of eastward flow, the transition is marked by the emergence of strong jets, initially as a single jet ($1\lesssim\xi\lesssim 10$) and later as multiple jets associated with potential vorticity staircases ($10\lesssim\xi$).
@@ 4 Linear Stability Analysis
Baroclinic instability in the Phillips model on the beta-plane is known to be asymmetric with respect to the zonal direction of the base flow (Pedlosky, 1987). This asymmetry is illustrated in figure 5, which shows the maximum growth rate as a function of $\delta$ and $\xi$ for both eastward and westward flows. The asymmetry appears in the intermediate stratification regime: westward flow becomes unstable for $\xi>1$, while eastward flow becomes unstable for $\xi>1/\delta$. The eastward flow in the gyres is hence expected to remain baroclinically stable in the intermediate stratification regime, while the westward flow of the gyres is baroclinically unstable. In the weak stratification regime, where surface-intensified flow in both directions is unstable, Sverdrup flow becomes barotropic according to potential vorticity homogenization theory, thus invalidating our initial hypothesis of a surface-intensified gyre.
@@
It follows an exponential function with a decay rate determined by the inertial layer thickness $\delta_{I}$. Additionally, a viscous sublayer is modeled by including another exponential term with a decay rate of $\delta_{P}$, bringing the velocity to zero near the western boundary. The velocity amplitude is determined by the inertial velocity $U_{I}=\psi_{Sv}/\delta_{I}$, ensuring that the gyre’s transport fully recirculates within this boundary layer profile. Further details can be found in Appendix C. The growth rate of the most unstable mode in this case is plotted as a function of $(\xi,\delta)$ in figure 5B and in figure 6B for the case $\delta=0.01$. An important result of this analysis is the existence of a stability island around $\xi=1$ for sufficiently small $\delta$. Near $\xi=1$, the western boundary stabilizes the meridional flow which should otherwise be baroclinically unstable (Spall, 2000). In figure 6B, we also plot the growth rate of the most unstable mode for the same velocity profile in a 1.5-layer quasi-geostrophic model, obtained by assuming $\psi_{2}=0$. This demonstrates that instabilities at $\xi<1$ are due to horizontal shear instabilities of the boundary layer, while instabilities at larger criticality parameters are genuinely baroclinic, as they are absent in the 1.5-layer model.
In figure 4 (right panels), we plot the wavelength of the most unstable mode for both the western meridional flow and the interior zonal flow, to compare this length scale with flow structures. For the interior flow, the most unstable mode has a length scale $2\pi L_{d}$, which appears to qualitatively match the size of the vortex rings. In contrast, for the western boundary current, the instability scale is initially much larger than the deformation radius, and only at values of $\xi>10$ approaches the order of the deformation radius. Snapshots of $\psi$ show the presence of large-scale basin modes (not shown), which may result from this instability at intermediate $\xi\sim 10$. At higher $\xi\sim 100$, baroclinic instability is supposedly of the same type as in the interior flow, as the instability scale becomes much smaller than the scale of the base profile, $\delta_{I}$. However, at this point the mean flow is no longer western intensified (figure 4A). Baroclinic instability itself may in fact be responsible for this erosion of Sverdrup flow, but a detailed analysis of the mean flow balance is beyond the scope of the present work.
@@
In the strong stratification regime ($\xi<1$), both models yield a similar flow pattern, resembling the Gyre Turbulence regime of Miller et al. (2024). This suggests that the dynamics are primarily governed by 1.5-layer dynamics, consistent with the linear computation predicting no baroclinic instability in this parameter range. As $\xi$ becomes larger than one, both models simulate a change of regime with the emergence of strong eastward jets. However, a notable difference occurs: recirculating westward jets appear in the 1.5-layer model but not in the two-layer model (see figure 7B and figure 7E). We argue that the mechanism for the apparition of the eastward jet is the same in both models, but that baroclinic instability of the westward jets is responsible for their absence in the two-layer model.
@@
So far, we have explained how turbulence may drive a change in the flow pattern within the domain bulk. We now discuss how these patterns may be connected with changes in the western boundary layer dynamics. The emergence of the eastward jet roughly coincides with the stabilization of the western boundary layer (Figure 5), and we hypothesize this stability to give rise to similar recirculation zones as observed in free-slip solutions (Ierley and Young, 1988). Excitation of different inertial recirculation modes, as outlined by Marshall and Marshall (1992), might also be relevant during the transition. In the asymptotic limits, the 1.5-layer solutions are characterised by tight, modon-like recirculation when $\xi\ll 1$ (mean flow not shown, but similar to left panels of figure 2) and by Fofonoff flow when $\xi\gg 1$, resembling inertial runaway (Sheremet et al. (1995), figure 7F). However, when $\xi\sim 1$, both a strong modon and a penetrating jet are present in the outflow region of the western boundary layer (figure 7E), and discussing the flow in terms of stationary inertial solutions only does not seem adequate. Although a mechanism setting the pattern of inertial recirculation is left to be identified, these runs are readily compared to the 2-layer simulations. Intense westward parts of inertial recirculation dissapear in the 2-layer runs (7B), which is attributed to zonal assymetry of baroclinic instability in surface intensified configurations.

### v2

# The Impact of Stratification on Surface-Intensified Eastward Jets in Turbulent Gyres

@@ Abstract
Eastward oceanic jets at mid-latitudes, such as the Gulf Stream or the Kuroshio, are important in oceanic circulation as they transport water masses from western boundary currents far into the open ocean. To isolate the role of density stratification on the formation of such jets in turbulent gyres, this study uses an idealized model with two layers of fluid stacked on top of each other. Our findings describe layer densities and depths that are favourable for jet formation, highlighting the importance of intensified flow in a thin surface layer.
@@ 1 Introduction
Surface-intensified eastward jets detaching from western boundaries are a prominent feature of mid-latitude oceans. Two iconic examples are the Gulf Stream and the Kuroshio, which appear as coherent, narrow, meandering ribbons surrounded by mesoscale oceanic rings in snapshots of surface kinetic energy (Chassignet and Xu, 2021). Satellite altimetry and state-estimate reconstructions show that these jets often maintain a coherent structure over long distances, despite the transition from coastal to open-ocean environments (Sánchez-Román et al., 2024; Chassignet and Marshall, 2008). Coherent eastward jets are also found across a wide range of models, from high-resolution primitive equation models (Uchida et al., 2022; Ajayi et al., 2020) to highly idealized quasi-geostrophic models (Holland, 1978; Berloff and McWilliams, 1999a). Despite their robust presence in observations and models, the understanding of the turbulent dynamics of eastward jets detaching from the western boundary remains incomplete. Here, we investigate the role of stratification in eastward jet formation using a simple two-layer quasi-geostrophic model of wind-driven ocean gyres.
When a western boundary current detaches it penetrates into the open ocean as an eastward jet, which tends to become unstable and shed eddies (Veronis, 1966; Holland, 1978). These eddies can reinforce and sharpen the jet as it flows away from the coast (Wardle and Marshall, 2000; Greatbatch et al., 2010; Dritschel and Scott, 2011). This has led numerous studies to neglect the connection of the jet to the western boundary current and instead investigate the interaction between turbulence and jets in periodic domains (Arbic and Flierl, 2003, 2004; Gallet and Ferrari, 2021), where jets may spontaneously emerge due to the planetary vorticity gradient (Rhines, 1975). However, it remains unclear how heterogeneity and the presence of a western boundary affect the dynamics of such freely-evolving flows (Nadiga and Straub, 2010; Grooms et al., 2013). Statistical mechanics may explain the spontaneous emergence of eastward jets in closed domains (Venaille and Bouchet, 2011), but fails to account for western intensification. Dynamical system theory can provide insights on the length of a viscous eastward jet detaching from a western boundary current (Simonnet, 2005), but is unfit to describe fully turbulent flow. To wit, there is no theoretical framework that can bridge the gap between turbulent eastward jets in the open ocean and western boundary currents.
In this study we aim to understand the formation of turbulent jets detaching from western boundaries by characterizing the necessary conditions for their emergence. We focus on the impact of stratification. Previous studies have shown that geostrophic turbulence can reinforce the jet in stratified models of ocean gyres (Sun et al., 2013), while the absence of stratification leads to its complete disintegration (Greatbatch and Nadiga, 2000; Fox-Kemper, 2005). Two stratification properties are known to influence turbulent flow dynamics in the ocean: the depth of the pycnocline and the amplitude of the associated density jump. A shallow pycnocline favors surface-intensified dynamics (Fu and Flierl, 1980; Smith and Vallis, 2001; Meunier et al., 2023). A stronger density jump increases the internal Rossby radius of deformation, a central horizontal length scale of geostrophic turbulence (Vallis, 2017). For instance, this length controls turbulent regime transitions from isolated vortices to jets separating regions of homogenized potential vorticity (Arbic and Flierl, 2003).
This article further explores the role of stratification by describing its impact on both basin-scale circulation and turbulent eastward jets. For this, we consider a two-layer quasi-geostrophic model of wind-driven ocean gyres which is the minimal model to describe both mesoscale turbulence and Sverdrup gyres. In this model, stratification is also characterized by two key parameters: the layer depth ratio, which translates to the relative depth of the pycnocline, and the internal Rossby radius of deformation. We seek to characterize transitions between flow regimes with and without jets when these two parameters are varied and interpret the results using a combination of linear stability analysis and geostrophic turbulence phenomenology.
@@
The structure of the article is as follows: in section 2, we describe the two-layer model and introduce the reduced parameter space which describes the vertical stratification properties. In section 3, we use numerical simulations to identify three distinct stratification regimes: strong, intermediate, and weak. We then shift our attention to the intermediate stratification regime. Section 4 presents a local linear stability analysis of surface-intensified gyres, paying particular attention to the distinct stability characteristics of the eastward and westward parts of the gyres, as well as the western boundary layer. In section 5, we compare a two-layer and a one-and-a-half-layer quasi-geostrophic simulation to identify the dynamics specifically linked to baroclinic instability. Finally, in section 6, we discuss the transition to a zonostrophic regime, which precedes the weak stratification regime characterized by barotropic gyres. We conclude in section 7 by outlining the minimal conditions required for the emergence and persistence of a coherent jet between the gyres.
@@ 2 Flow Model
Wind Stress $\tau_{0}$ $50\times 10^{-5}$ m2 s-2, $16\times 10^{-5}$ m2 s-2, $1\times 10^{-5}$ m2 s-2
Sverdrup Velocity scale $U_{Sv}$ $0.09$ m2 s-1
@@
with $\psi_{i}$ the streamfunction and $\mathbf{u}_{i}=(u_{i},v_{i})=(-\partial_{y}\psi_{i},\partial_{x}\psi_{i})$ the velocity in each layer. We solve this system of equations on a square domain with $(x,y)\in[0,L]^{2}$. Since our study focuses primarily on regimes with surface-intensified turbulent flow, we consistently neglect linear drag which is often thought to represent bottom friction in the bulk of the domain. We instead choose to dissipate energy along the domain boundary by applying lateral no-slip boundary conditions (Miller et al., 2024).
with $H_{i}$ being the layer thicknesses, $g^{\prime}$ the reduced gravity between both layers and $f_{0}$ the Coriolis parameter. In addition to $\delta$, the problem admits three non-dimensional parameters:
$\tilde{\nu}=\frac{\nu}{U_{Sv}L},\quad\tilde{\beta}=\frac{\beta L^{2}}{U_{Sv}},\quad\xi=\frac{U_{Sv}}{\beta L_{d}^{2}}$ (6)
with
$U_{Sv}=\frac{8\pi\tau_{0}}{H_{1}L\beta}$ (7)
being the velocity scale of the Sverdrup flow. For all simulations, we set $L=4000$ km, $\beta=1.7\times 10^{-11}$ m-1 s-1, $\nu=10$ m2 s-1, and $U_{Sv}=0.09$ m s-1. The parameters $\tilde{\nu}$ and $\tilde{\beta}$ are thus kept constant at $\tilde{\beta}\simeq 3000$, $\tilde{\nu}\simeq 10^{-5}$. This corresponds to a thickness of the inertial boundary layer $\delta_{I}=L/\sqrt{\tilde{\beta}}=72$ km (Charney, 1955) and a viscous sublayer of $\delta_{P}=L\sqrt{\tilde{\nu}/\sqrt{\tilde{\beta}}}=3$ km (Pedlosky, 1987; Ierley and Ruehr, 1986).
In this study, we focus on the dynamical regimes that appear upon varying the two parameters $\xi$ and $\delta$. $\delta$ is a non-dimensional measure of surface-intensification of the stratification. In two-layer models of planetary gyres, $\delta$ may be interpreted as the position of the pycnocline and is often placed at $\delta=0.1-0.2$ (Flierl, 1978). However, as shown in Smith and Vallis (2001), the scenario $\delta\ll 1$ is an interesting asymptotic limit because it corresponds to a situation where barotropization is halted at the first baroclinic mode, contrary to the classical picture where energy cascades to the barotropic mode (Salmon, 1998). In order to investigate this limit, we will consider the range $0.01\leq\delta\leq 0.5$ and focus mostly on $\delta=0.01$. Although this value is probably beyond the realistic value observed in the ocean, the dynamical regimes were easier to entangle for such a small value of $\delta$.
The parameter $\xi$ is a non-dimensional deformation radius sometimes called the criticality parameter, related to baroclinic instability of the two-layer model (Stone, 1978). We recall here the interpretation of Held and Larichev (1996) and Jansen and Ferrari (2012) adapted to our configuration. Suppose there is a zonal Sverdrup flow of magnitude $U_{Sv}$ confined to the upper layer. The associated potential vorticity gradient in the upper and lower layer is
$\frac{\partial Q_{1}}{\partial y}=\beta+\frac{(1-\delta)U_{Sv}}{L_{d}^{2}}\,,\quad\textrm{and}\quad\frac{\partial Q_{2}}{\partial y}=\beta-\frac{\delta U_{Sv}}{L_{d}^{2}}\,.$ (8)
For baroclinic instability to occur, these gradients have to be of opposite sign (Pedlosky, 1987). For westward flow ($U_{Sv}<0$) the gradient of the lower layer is always positive, and the condition for baroclinic instability is reached when $\partial Q_{1}/\partial y<0$. In the limit $\delta\ll 1$, this corresponds to $\xi>1$. For eastward flow ($U_{Sv}>0$) the gradient of the upper layer is always positive and the condition for instability is given by $\partial Q_{2}/\partial y<0$, which corresponds to $\delta\xi>1$. $\xi$ therefore serves as a proxy for the baroclinic stability of the large-scale flow, where $\xi<1$ corresponds to stable flows. Furthermore, there is a second interpretation of $\xi$ in ocean gyres because we can write
$\xi=\frac{\delta_{I}^{2}}{L_{d}^{2}}\,,$ (9)
Then, $\xi$ is a non-dimensional measure of the width of the inertial boundary layer relative to the deformation radius.
In order to reach asymptotic behaviour, we vary $\xi$ from $10^{-2}$ to $10^{2}$ for $\delta=0.5$, from $1$ to $5.6$ for $\delta=1/6$ and from $10^{-1}$ to $10^{3}$ for $\delta=0.01$. This corresponds to changing $L_{d}$ from $737$ km to $7$ km, $72$ km to $13$ km and $233$ km to $2$ km, respectively. In our numerical simulations, we change $L_{d}$ by setting $g^{\prime}$, which controls $\xi$ but leaves all other non-dimensional variables unchanged. A summary of the dimensional variables of the simulations is given in table 1.
To numerically solve equations (1) - (2) the code qgw was used. It uses classical finite differences discretization of the equations with an Arakawa formulation of the Jacobian and a direct spectral inversion (after projection on the vertical modes) to obtain the streamfunction $\psi$ from the potential vorticity $q$. Time stepping is done with an adaptive 2nd order Adams-Bashforth technique, where the time step is determined by a standard CFL-condition. Unless specified, all runs are performed with 4096$\times$4096 grid points on the horizontal, corresponding to a resolution of $1$ km, ensuring all relevant flow scales of the problem are well resolved. We paid attention to run the simulations for at least two gyre turnover times $T_{Gyre}=L/U_{Sv}$ after the spinup was completed. Last, the no-slip boundary condition is implemented as in Miller et al. (2024).
@@ 3.1 Reference Run
Figure 1: Upper layer (top row) and lower layer (bottom row) for the reference run at $\delta=0.01$, $\xi=3.3$. (A, D) Flow speed $|\mathbf{u}_{i}|$ normalized with inertial velocity $U_{I}=\psi_{Sv}/\delta_{I}$, (B, E) potential vorticity normalised by planetary vorticity and (C, F) time-mean streamfunction normalised by Sverdrup scaling.
@@ 3.2 Parameter Space
The focus of this manuscript lies on the intermediate stratification regime, due to the appearance of eastward jets in gyres such as presented in the reference run. This regime is characterised by $1<\xi<1/\delta$ and is hence accessible only when $\delta\ll 1$. We therefore focus on the runs at $\delta=0.01$, limiting our description to the upper layer as the dynamics remain mostly surface-intensified (figure 3). The surface flow patterns are displayed in figure 4. As $\xi$ is increased, a continuous transition occurs from western-intensified gyres to a state that restores east-west symmetry, with a decrease in total transport of the gyres. Coincidentally, the dominant turbulent flow features change. The strong coherent vortices which dominate the flow in the strong stratification regime disappear, and instead a sea of eddies of characteristic size $L_{d}$ develops in the regions of westward flow. In the regions of eastward flow, the transition is marked by the emergence of strong jets, initially as a single jet ($1\lesssim\xi\lesssim 10$) and later as multiple jets associated with potential vorticity staircases ($10\lesssim\xi$).
@@ 4 Linear Stability Analysis
Baroclinic instability in the Phillips model on the beta-plane is known to be asymmetric with respect to the zonal direction of the mean flow (Pedlosky, 1987). This asymmetry is illustrated in figure 5, which shows the maximum growth rate as a function of $\delta$ and $\xi$ for both eastward and westward Sverdrup flows. The asymmetry appears in the intermediate stratification regime: westward Sverdrup flow becomes unstable for $\xi>1$, while eastward Sverdrup flow becomes unstable for $\xi>1/\delta$. In the weak stratification regime, where surface-intensified flow in both directions is unstable, Sverdrup flow becomes barotropic according to potential vorticity homogenization theory, thus invalidating our initial hypothesis of a surface-intensified gyre.
@@
It follows an exponential function with a decay rate determined by the inertial layer thickness $\delta_{I}$. Additionally, a viscous sublayer is modeled by including another exponential term with a decay rate of $\delta_{P}$, bringing the velocity to zero near the western boundary. The velocity amplitude is determined by the inertial velocity $U_{I}=\psi_{Sv}/\delta_{I}$, ensuring that the gyre’s transport fully recirculates within this boundary layer profile. Further details can be found in Appendix C. The growth rate of the most unstable mode in this case is plotted as a function of $(\xi,\delta)$ in figure 5B and in figure 6B for the case $\delta=0.01$. An important result of this analysis is the existence of a stability island around $\xi=1$ for sufficiently small $\delta$. Near $\xi=1$, we observe that a meridional flow with characteristic width $\delta_{I}\simeq L_{d}$ is stable. This adds to the work of Spall (2000) who showed that a uniform meridional flow is necessarily baroclinically unstable. In figure 6B, we also plot the growth rate of the most unstable mode for the same velocity profile in a 1.5-layer quasi-geostrophic model, obtained by assuming $\psi_{2}=0$. This demonstrates that instabilities at $\xi<1$ are due to horizontal shear instabilities of the boundary layer, while instabilities at larger criticality parameters are genuinely baroclinic, as they are absent in the 1.5-layer model.
In figure 4 (right panels), we plot the wavelength of the most unstable mode for both the western meridional flow and the zonal Sverdrup flow, to compare this length scale with flow structures. For the Sverdrup flow, the most unstable mode has a length scale $2\pi L_{d}$, which appears to qualitatively match the size of the vortex rings. In contrast, for the western boundary current, the instability scale is initially much larger than the deformation radius, and only at values of $\xi>10$ approaches the order of the deformation radius. Snapshots of $\psi$ show the presence of large-scale basin modes (not shown), which may result from this instability at intermediate $\xi\sim 10$. At higher $\xi\sim 100$, baroclinic instability is supposedly of the same type as in the Sverdrup flow, as the instability scale becomes much smaller than the scale of the mean profile, $\delta_{I}$. However, at this point the mean flow is no longer western intensified (figure 4A). Baroclinic instability itself may in fact be responsible for this erosion of Sverdrup flow, but a detailed analysis of the mean flow balance is beyond the scope of the present work.
@@
In the strong stratification regime ($\xi<1$), both models yield a similar flow pattern, resembling the Gyre Turbulence regime of Miller et al. (2024). This suggests that the dynamics are primarily governed by 1.5-layer dynamics, consistent with the linear computation predicting no baroclinic instability in this parameter range. As $\xi$ becomes larger than one, both models simulate a change of regime with the emergence of strong eastward jets. However, a notable difference occurs: recirculating westward jets appear in the 1.5-layer model but not in the two-layer model (see figure 7B and figure 7E). We attribute this difference to the zonal asymmetry of baroclinic instability in the intermediate stratification regime. If westward Sverdrup flow is baroclinically unstable then westwards jets will be unstable, too, explaining their disappearance in the 2-layer model. However, we argue that the mechanism for jet creation is the same in both models.
@@
So far, we have explained how turbulence may drive a change in the flow pattern within the domain bulk. We now discuss how these patterns may be connected with changes in the western boundary layer dynamics. The emergence of the eastward jet roughly coincides with the stabilization of the western boundary layer (Figure 5), and we hypothesize this stability to give rise to similar recirculation zones as observed in free-slip solutions (Ierley and Young, 1988). Excitation of different inertial recirculation modes, as outlined by Marshall and Marshall (1992), might also be relevant during the transition. In the asymptotic limits, the 1.5-layer solutions are characterised by a pair of contra-rotating vortices near the western boundary when $\xi\ll 1$ (mean flow not shown, but similar to left panels of figure 2) and by Fofonoff flow when $\xi\gg 1$, resembling inertial runaway (Sheremet et al. (1995), figure 7F). However, when $\xi\sim 1$, both a strong modon and a penetrating jet are present in the outflow region of the western boundary layer (figure 7E), and discussing the flow in terms of stationary inertial solutions only does not seem adequate. Although a mechanism setting the pattern of inertial recirculation is left to be identified, these runs are readily compared to the 2-layer simulations. Intense westward parts of inertial recirculation dissapear in the 2-layer runs (7B), which is attributed to zonal assymetry of baroclinic instability in surface intensified configurations.